# Learning to Project for Cross-Task Knowledge Distillation

### v1

@@ Abstract
Traditional knowledge distillation (KD) relies on a proficient teacher trained on the target task, which is not always available. In this setting, cross-task distillation can be used, enabling the use of any teacher model trained on a different task. However, many KD methods prove ineffective when applied to this cross-task setting. To address this limitation, we propose a simple modification: the use of an inverted projection. We show that this drop-in replacement for a standard projector is effective by learning to disregard any task-specific features which might degrade the student’s performance. We find that this simple modification is sufficient for extending many KD methods to the cross-task setting, where the teacher and student tasks can be very different. In doing so, we obtain up to a 1.9% improvement in the cross-task setting compared to the traditional projection, at no additional cost. Our method can obtain significant performance improvements (up to 7%) when using even a randomly-initialised teacher on various tasks such as depth estimation, image translation, and semantic segmentation, despite the lack of any learned knowledge to transfer. To provide conceptual and analytical insights into this result, we show that using an inverted projection allows the distillation loss to be decomposed into a knowledge transfer and a spectral regularisation component. Through this analysis we are additionally able to propose a novel regularisation loss that allows teacher-free distillation, enabling performance improvements of up to 8.57% on ImageNet with no additional training costs.
(a) Traditional same-task knowledge distillation that projects to the teacher feature space.
(b) Cross-task knowledge distillation using our inverted projection.
Figure 1: Our cross-task knowledge distillation pipeline, where a student model is trained on a target task with the aid of a frozen teacher that is pretrained on a different task. Compared to standard same-task feature distillation (fig. 0(a)), our cross-task approach uses an inverted projector (fig. 0(b)) which is able to discard irrelevant task-specific features from the different-task teacher. The loss comprises a feature distillation loss $\mathcal{L}_{distill}$ that matches the student features with the projected teacher features, and a task-specific supervised loss $\mathcal{L}_{task}$ applied only to the student model’s output for the target task. Refer to caption denotes trainable weights, while Refer to caption denotes weights that are frozen throughout training.
@@ 1 Introduction
Knowledge distillation (KD) has emerged as a very effective tool for training small and efficient models [70, 52, 21, 5, 49, 44]. It leverages the pre-trained knowledge of a much larger (teacher) model to guide and enhance the training process of a significantly smaller (student) model. Since its inception, KD has been applied to a wide variety of tasks in the computer vision [10], audio [14] and language [64] domains, enabling the deployment of models across many embedded devices.
However, existing approaches for KD are often limited to the cases where the teacher shares the same task with the student [31, 11, 29, 70, 9, 10]. This is very restrictive since there are many applications where there is simply no suitable pretrained teacher available due to, for example, the lack of any large annotated training data. This problem commonly arises for tasks that require expensive human annotation [46], such as in robotics [33], or where the collection of data is prohibitive for other reasons, such as in the medical [63, 38] and aerial domains [74, 23, 36]. In these cases, it is not possible to train a suitable teacher for the target task, therefore we propose a cross-task knowledge distillation. In the cross-task KD setup, a teacher model trained for a different task can be used to improve the student performance. This setting is well-suited for tasks lacking a task-specific pretrained teacher, as it allows for any other off-the-shelf pretrained model to be used to improve the student model performance instead. It is also increasingly relevant as the compute and data costs to train large models increases. Similarly, data labelling may be cheaper for one task than another, e.g. training a model using a cheaply-labelled auxiliary task is very common in active learning [4, 25, 30] and federated learning [1].
@@ 2 Related Work
Knowledge distillation (KD) is a technique that involves transferring the knowledge from a large teacher model to a smaller student model, aiming to improve the performance of the student on the target task. It has become increasingly popular in recent years for the deployment of models on resource constrained devices, such as mobile phones, and has been applied in image classification [29, 9], semantic segmentation [42], video object segmentation [52], and natural language processing [65]. Existing literature has extensively explored various distillation pipelines [45, 39, 60] along with both empirical and theoretically motivated loss formulations [29, 82, 9, 51] that can facilitate the knowledge transfer process. However, these conventional methods still predominantly focus on same-task distillation [70, 10], wherein the student and teacher models are trained on the same task. There are many applications where there are no pre-trained off-the-shelf teacher models available, thus motivating the need to perform cross-task distillation. Some prior works have pursued cross-task distillation both as a generalisation of the knowledge transfer occurring in traditional knowledge distillation [78] and because of the observation that some tasks will naturally tend to share information [79, 76]. CrossDistil [76] was one of the first to partially explore this new setting by introducing a quadruplet loss, calibration term, and an error correction mechanism, however knowledge was distilled between the task-specific decoder heads of a multi-task model with shared encoder weights, rather than between two fully-separate models. ProC-KD [40] transfer local-level object knowledge for large-scale cross-task distillation, while [78] construct a relational embedding for the loss. [79] perform cross-task KD to augment text-to-image generation using image semantic encoders, but the proposed method is tightly coupled with the model architecture at each stage of the pipeline. In contrast to these works, we propose a very simple extension to the typical feature distillation pipeline that enables the distillation of knowledge cross-task across a wide range of settings.
@@
Transfer learning and domain adaptation are widely studied areas in machine learning that leverage the knowledge acquired by a pretrained model on one task to enhance the performance on a different, yet related task [85]. This paradigm has demonstrated significant success in various fields, especially in computer vision [81, 20] and natural language processing [58, 68], by reducing the training time and data requirements in the target domain. Most existing transfer learning or domain adaptation algorithms attempt to align the feature representations across the two domains. This can be achieved by minimising some statistical discrepancy between the two spaces [27] or introducing additional adversarial losses [17]. More recent works [13] have shown a spectral divide between the domain-specific and domain-agnostic features. This is where the large singular values of the features can generalise across domains, whereas the small singular values are domain-specific. This observation has led to follow-up works [13, 12, 55] by proposing spectral alignment and normalisation techniques. We take a similar approach for transferring knowledge between different tasks, but in the context of knowledge distillation, where an additional capacity gap between the source and target task models exists. This work also enables a more concrete bridge between the field of transfer learning and knowledge distillation.
@@
Multi-task learning. There are many cases where jointly training on multiple tasks or modalities can improve not only the generality of models [56] but also the single-task performance. For example, monocular depth estimation has been shown to share knowledge with other tasks, such as semantic segmentation [59, 34, 3, 75, 73, 2, 37]. Intuitively, this follows for other task pairs; for instance, both semantic segmentation and classification target the semantics within an image. Unfortunately, multitask models are often too large and expensive to run on resource-constrained devices [35]. Additionally, jointly learning multiple tasks with a small model can degrade the downstream performance, as additional tasks or objectives can conflict with the target task when there is insufficient capacity in the student to optimise for both [22]. This further motivates our approach to training efficient models for a single task by leveraging cross-task information.
@@ 3 Method
This section describes our proposed inverted projection and our pipeline for cross-task distillation. We first describe the cross-task knowledge distillation setting and detail why the typical approaches for same-task KD which use the traditional projector are not suited to it, particularly when there is a significant difference between the teacher and student tasks. We then present our inverted projection and analyse its dynamics to help understand the underlying principles of cross-task distillation. Finally, we use this analysis to derive a straightforward drop-in teacher-free distillation loss that provides some of the benefits of knowledge distillation without the use of any teacher model.
@@ 3.1 Cross-task Feature Distillation
Cross-task distillation is motivated by the intuitive and demonstrated overlap in useful information between different tasks (see section 2). We use feature-space distillation, which aims to align the feature spaces of a student model and a teacher model. To do this, a learnable projection is used to map the features from one model into the feature space of the other. However, in the cross-task case, where the teacher model has been trained for a significantly different task to the student model’s target task, there are specific issues to contend with that traditional KD methods do not address.
We introduce a novel inverted projection that is well-suited to the cross-task setting, in contrast to the traditional projection [62, 70] which is better suited to the same-task setting. In section 3.2, we provide the motivation for our method, and in section 3.3 we describe our pipeline, problem setup, and loss function.
@@ 3.2 Importance of Feature Projection
In the traditional same-task setting, the teacher model is already pre-trained for the student’s target task, so it is desirable for the student to match features as closely as possible with those produced by the teacher. In this case, the task-specific knowledge is helpful in improving the student performance. However, for the cross-task setting, the teacher is trained on a different task to the student. As detailed in section 2, there is likely at least some shared knowledge between different tasks. The issue in the cross-task knowledge distillation setting is how to extract only the task-agnostic knowledge and the knowledge shared between tasks, all while ignoring the irrelevant features produced by teacher. This last point is especially important for smaller student models as they do not have the capacity to effectively learn the union of two very different feature spaces.
A projection layer is often used in knowledge distillation to match the student’s feature dimensions with those of the teacher [62, 70]. Although recent works have highlighted the importance of the projector in improving the efficacy of same-task distillation [15, 50, 48], they have proven ineffective when there is cross-task information present. We propose a modification of the projection in which we instead map from the teacher space onto the student space. We describe this as inverting the projector, and we find that it enables the suppression of irrelevant (task-specific) features. We show that this inverted projector can effectively discard these irrelevant features if needed. If this were to be used in the traditional same-task setting, the discarding of these features would be detrimental, but in the cross-task setting, it is actively desirable.
Table 1: Cross-task distillation to a depth estimation student model using similar ( normal-$\blacksquare$) and dissimilar ( normal-$\blacksquare$) teacher tasks, showing the increasing effect of our inverted projection as similarity between teacher and student tasks decreases. We use our inverted projector with four different KD methods to show its general applicability. The inverted projector outperforms traditional projections in the cross-task case for which it is designed, but always produces a performance improvement over the baseline (no distillation) regardless of the teacher task. normal-$\blacksquare$ normal-$\blacksquare$ normal-$\blacksquare$ normal-$\blacksquare$ colour map denotes decreasing student-teacher task similarity. Results on NYUv2 [66] using an EfficientNet-B0 [69]-based student.
Teacher task normal-$\longrightarrow$ $\blacksquare$ Depth $\blacksquare$ Instance Seg. $\blacksquare$ Classification $\blacksquare$ Random
@@
FitNets [62] ICLR 2015 Traditional 0.868 0.117 0.406 0.855 0.122 0.425 0.845 0.125 0.439 0.828 0.134 0.455
@@
AT [80] ICLR 2017 Traditional 0.856 0.122 0.426 0.852 0.123 0.431 0.850 0.125 0.433 0.857 0.121 0.428
@@
PKT [54] ECCV 2018 Traditional 0.854 0.122 0.429 0.857 0.123 0.427 0.851 0.124 0.432 0.856 0.123 0.429
@@
Ensemble [15] NeurIPS 2022 Traditional 0.861 0.119 0.416 0.856 0.122 0.425 0.852 0.124 0.431 0.835 0.128 0.446

### v2

# Imperial College London South Kensington, UK Learning to Project for Cross-Task Knowledge Distillation

@@ Abstract
Traditional knowledge distillation (KD) relies on a proficient teacher trained on the target task, which is not always available. In this setting, cross-task distillation can be used, enabling the use of any teacher model trained on a different task. However, many KD methods prove ineffective when applied to this cross-task setting. To address this limitation, we propose a simple modification: the use of an inverted projection. We show that this drop-in replacement for a standard projector is effective by learning to disregard any task-specific features which might degrade the student’s performance. We find that this simple modification is sufficient for extending many KD methods to the cross-task setting, where the teacher and student tasks can be very different. In doing so, we obtain up to a 1.9% improvement in the cross-task setting compared to the traditional projection, at no additional cost. Our method can obtain significant performance improvements (up to 7%) when using even a randomly-initialised teacher on various tasks such as depth estimation, image translation, and semantic segmentation, despite the lack of any learned knowledge to transfer. To provide conceptual and analytical insights into this result, we show that using an inverted projection allows the distillation loss to be decomposed into a knowledge transfer and a spectral regularisation component. Through this analysis we are additionally able to propose a novel regularisation loss that allows teacher-free distillation, enabling performance improvements of up to 2.3% on ImageNet with no additional training costs.
The authors contributed equally to this paper
@@ 1 Introduction
Knowledge distillation (KD) has emerged as a very effective tool for training small and efficient models [Tian et al.(2019)Tian, Krishnan, and Isola, Miles et al.(2023)Miles, Yucel, Manganelli, and Saa-Garriga, Fang et al.(2021)Fang, Song, Wang, Shen, Wang, and Song, Bhardwaj et al.(2019)Bhardwaj, Suda, and Marculescu, Miles and Mikolajczyk(2020), Lopez-Paz et al.(2016)Lopez-Paz, Bottou, Schölkopf, and Vapnik]. It leverages the pre-trained knowledge of a much larger (teacher) model to guide and enhance the training process of a significantly smaller (student) model. Since its inception, KD has been applied to a wide variety of tasks in the computer vision [Chen et al.(2021a)Chen, Liu, Zhao, and Jia], audio [Chen et al.(2021b)Chen, Xian, Koepke, Shan, and Akata] and language [Sanh et al.(2019a)Sanh, Debut, Chaumond, and Wolf] domains, enabling the deployment of models across many embedded devices.
However, existing approaches for KD are often limited to the cases where the teacher shares the same task with the student [Huang et al.(2022)Huang, You, Wang, Qian, and Xu, Chen et al.(2022a)Chen, Cao, Zhong, Zhang, Gao, and Tao, Hinton et al.(2015)Hinton, Vinyals, and Dean, Tian et al.(2019)Tian, Krishnan, and Isola, Chen et al.(2020)Chen, Wang, Gan, Liu, Henao, and Carin, Chen et al.(2021a)Chen, Liu, Zhao, and Jia]. This is very restrictive since there are many applications where there is simply no suitable pretrained teacher available due to, for example, the lack of any large annotated training data. This problem commonly arises for tasks that require expensive human annotation [Mccormac et al.(2017)Mccormac, Handa, Leutenegger, and Davison], such as in robotics [James et al.(2019)James, Wohlhart, Kalakrishnan, Kalashnikov, Irpan, Ibarz, Levine, Hadsell, and Bousmalis], or where the collection of data is prohibitive for other reasons, such as in the medical [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox, Komorowski et al.(2023)Komorowski, Baniecki, and Biecek] and aerial domains [Wang et al.(2020b)Wang, Zhu, Wang, Hu, Qiu, Wang, Hu, Kapoor, and Scherer, Fonder and Van Droogenbroeck(2019), Kolbeinsson and Mikolajczyk(2023)]. In these cases, it is not possible to train a suitable teacher for the target task, therefore we propose a cross-task knowledge distillation. In the cross-task KD setup, a teacher model trained for a different task can be used to improve the student performance. This setting is well-suited for tasks lacking a task-specific pretrained teacher, as it allows for any other off-the-shelf pretrained model to be used to improve the student model performance instead. It is also increasingly relevant as the compute and data costs to train large models increases. Similarly, data labelling may be cheaper for one task than another, e.g. training a model using a cheaply-labelled auxiliary task is very common in active learning [Baldridge and Osborne(2004), Gao and Saar-Tsechansky(2020), Huang et al.(2017)Huang, Chen, Mu, and Zhou] and federated learning [Ahn et al.(2022)Ahn, Kim, Koh, and Li].
@@ 2 Related Work
Knowledge distillation (KD) is a technique that involves transferring the knowledge from a large teacher model to a smaller student model, aiming to improve the performance of the student on the target task. It has become increasingly popular in recent years for the deployment of models on resource constrained devices, such as mobile phones, and has been applied in image classification [Hinton et al.(2015)Hinton, Vinyals, and Dean, Chen et al.(2020)Chen, Wang, Gan, Liu, Henao, and Carin], semantic segmentation [Liu et al.(2019)Liu, Chen, Liu, Qin, Luo, and Wang], video object segmentation [Miles et al.(2023)Miles, Yucel, Manganelli, and Saa-Garriga], and natural language processing [Sanh et al.(2019b)Sanh, Debut, Chaumond, and Wolf]. Existing literature has extensively explored various distillation pipelines [Malinin et al.(2020)Malinin, Mlodozeniec, and Gales, Le et al.(2020)Le, Vo, and Thoa, Ren et al.(2022)Ren, Gao, Hua, Xue, Tian, He, and Zhao] along with both empirical and theoretically motivated loss formulations [Hinton et al.(2015)Hinton, Vinyals, and Dean, Zhao et al.(2022)Zhao, Song, and Qiu, Chen et al.(2020)Chen, Wang, Gan, Liu, Henao, and Carin, Miles et al.(2022)Miles, Rodriguez, and Mikolajczyk] that can facilitate the knowledge transfer process. However, these conventional methods still predominantly focus on same-task distillation [Tian et al.(2019)Tian, Krishnan, and Isola, Chen et al.(2021a)Chen, Liu, Zhao, and Jia], wherein the student and teacher models are trained on the same task. There are many applications where there are no pre-trained off-the-shelf teacher models available, thus motivating the need to perform cross-task distillation. Some prior works have pursued cross-task distillation both as a generalisation of the knowledge transfer occurring in traditional knowledge distillation [Ye et al.(2020)Ye, Lu, and Zhan] and because of the observation that some tasks will naturally tend to share information [Yuan and Peng(2020), Yang et al.(2022)Yang, Pan, Gao, Jiang, Liu, and Chen]. CrossDistil [Yang et al.(2022)Yang, Pan, Gao, Jiang, Liu, and Chen] was one of the first to partially explore this new setting by introducing a quadruplet loss, calibration term, and an error correction mechanism, however knowledge was distilled between the task-specific decoder heads of a multi-task model with shared encoder weights, rather than between two fully-separate models. ProC-KD [Li et al.(2022)Li, Wu, Han, and Tian] transfer local-level object knowledge for large-scale cross-task distillation, while [Ye et al.(2020)Ye, Lu, and Zhan] construct a relational embedding for the loss. [Yuan and Peng(2020)] perform cross-task KD to augment text-to-image generation using image semantic encoders, but the proposed method is tightly coupled with the model architecture at each stage of the pipeline. In contrast to these works, we propose a very simple extension to the typical feature distillation pipeline that enables the distillation of knowledge cross-task across a wide range of settings.
@@
Transfer learning and domain adaptation are widely studied areas in machine learning that leverage the knowledge acquired by a pretrained model on one task to enhance the performance on a different, yet related task [Zhuang et al.(2020)Zhuang, Qi, Duan, Xi, Zhu, Zhu, Xiong, and He]. This paradigm has demonstrated significant success in various fields, especially in computer vision [Zamir et al.(2018)Zamir, Sax, Shen, Guibas, Malik, and Savarese, Evci et al.(2022)Evci, Dumoulin, Larochelle, and Mozer] and natural language processing [Raffel et al.(2020)Raffel, Shazeer, Roberts, Lee, Narang, Matena, Zhou, Li, and Liu, Sung et al.(2022)Sung, Cho, and Bansal], by reducing the training time and data requirements in the target domain. Most existing transfer learning or domain adaptation algorithms attempt to align the feature representations across the two domains. This can be achieved by minimising some statistical discrepancy between the two spaces [Greenfeld and Shalit(2019)] or introducing additional adversarial losses [Deng et al.(2021)Deng, Zhang, Vodrahalli, Kawaguchi, and Zou]. More recent works [Chen et al.(2019b)Chen, Wang, Long, and Wang] have shown a spectral divide between the domain-specific and domain-agnostic features. This is where the large singular values of the features can generalise across domains, whereas the small singular values are domain-specific. This observation has led to follow-up works [Chen et al.(2019b)Chen, Wang, Long, and Wang, Chen et al.(2019a)Chen, Wang, Fu, Long, and Wang, Raab et al.(2020)Raab, Väth, Meier, and Schleif] by proposing spectral alignment and normalisation techniques. We take a similar approach for transferring knowledge between different tasks, but in the context of knowledge distillation, where an additional capacity gap between the source and target task models exists. This work also enables a more concrete bridge between the field of transfer learning and knowledge distillation.
@@
Multi-task learning. There are many cases where jointly training on multiple tasks or modalities can improve not only the generality of models [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, Krueger, and Sutskever] but also the single-task performance. For example, monocular depth estimation has been shown to share knowledge with other tasks, such as semantic segmentation [Ramamonjisoa and Lepetit(2019), Jiao et al.(2018)Jiao, Cao, Song, and Lau, Bai et al.(2019)Bai, Fan, Pan, and Chen, Xing et al.(2022)Xing, Shen, Ho, and Tzes, Wang et al.(2020a)Wang, Zhang, Wang, Lin, and Lu, Auty and Mikolajczyk(2022), Kolbeinsson and Mikolajczyk(2024)]. Intuitively, this follows for other task pairs; for instance, both semantic segmentation and classification target the semantics within an image. Unfortunately, multitask models are often too large and expensive to run on resource-constrained devices [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick]. Additionally, jointly learning multiple tasks with a small model can degrade the downstream performance, as additional tasks or objectives can conflict with the target task when there is insufficient capacity in the student to optimise for both [Fifty et al.(2021)Fifty, Amid, Zhao, Yu, Anil, and Finn].
@@ 3.1 Cross-task Feature Distillation
Cross-task distillation is motivated by the intuitive and demonstrated overlap in useful information between different tasks (see section 2). We use feature-space distillation, which aims to align the feature spaces of a student model and a teacher model. To do this, a learnable projection is used to map the features from one model into the feature space of the other. However, in the cross-task case, where the teacher model has been trained for a significantly different task to the student model’s target task, there are specific issues to contend with that traditional KD methods do not address. We introduce a novel inverted projection that is well-suited to the cross-task setting, in contrast to the traditional projection [Romero et al.(2015b)Romero, Ballas, Kahou, Chassang, Gatta, and Bengio, Tian et al.(2019)Tian, Krishnan, and Isola] which is better suited to the same-task setting.
@@ 3.2 Importance of Feature Projection
In the traditional same-task setting, the teacher model is already pre-trained for the student’s target task, so it is desirable for the student to match features as closely as possible with those produced by the teacher. In this case, the task-specific knowledge is helpful in improving the student performance. However, for the cross-task setting, the teacher is trained on a different task to the student. As detailed in section 2, there is likely at least some shared knowledge between different tasks. The issue in the cross-task knowledge distillation setting is how to extract only the task-agnostic knowledge and the knowledge shared between tasks, all while ignoring the irrelevant features produced by teacher. This last point is especially important for smaller student models as they do not have the capacity to effectively learn the union of two very different feature spaces. A projection layer is often used in knowledge distillation to match the student’s feature dimensions with those of the teacher [Romero et al.(2015b)Romero, Ballas, Kahou, Chassang, Gatta, and Bengio, Tian et al.(2019)Tian, Krishnan, and Isola]. Although recent works have highlighted the importance of the projector in improving the efficacy of same-task distillation [Chen et al.(2022b)Chen, Wang, Liu, Xu, de Hoog, and Huang, Miles and Mikolajczyk(2024), Miles et al.(2024)Miles, Elezi, and Deng], they have proven ineffective when there is cross-task information present. We propose a modification of the projection in which we instead map from the teacher space onto the student space. We describe this as inverting the projector, and we find that it enables the suppression of irrelevant (task-specific) features. We show that this inverted projector can effectively discard these irrelevant features if needed. If this were to be used in the traditional same-task setting, the discarding of these features would be detrimental, but in the cross-task setting, it is actively desirable.
Teacher task $\longrightarrow$ $\blacksquare$ Depth $\blacksquare$ Instance Seg. $\blacksquare$ Classification $\blacksquare$ Random
@@
FitNets [Romero et al.(2015b)Romero, Ballas, Kahou, Chassang, Gatta, and Bengio] ICLR 2015 Traditional 0.868 0.117 0.406 0.855 0.122 0.425 0.845 0.125 0.439 0.828 0.134 0.455
@@
AT [Zagoruyko and Komodakis(2019)] ICLR 2017 Traditional 0.856 0.122 0.426 0.852 0.123 0.431 0.850 0.125 0.433 0.857 0.121 0.428
@@
PKT [Passalis and Tefas(2018)] ECCV 2018 Traditional 0.854 0.122 0.429 0.857 0.123 0.427 0.851 0.124 0.432 0.856 0.123 0.429
@@
Ensemble [Chen et al.(2022b)Chen, Wang, Liu, Xu, de Hoog, and Huang] NeurIPS 2022 Traditional 0.861 0.119 0.416 0.856 0.122 0.425 0.852 0.124 0.431 0.835 0.128 0.446
@@
Table 1: Cross-task distillation to a depth estimation student model using similar ( $\blacksquare$) and dissimilar ( $\blacksquare$) teacher tasks, showing the increasing effect of our inverted projection as similarity between teacher and student tasks decreases. We use our inverted projector with four different KD methods to show its general applicability. The inverted projector outperforms traditional projections in the cross-task case for which it is designed, but always produces a performance improvement over the baseline (no distillation) regardless of the teacher task. $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ colour map denotes decreasing student-teacher task similarity.